OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (3x - 7) units and the length of side ON is (x + 13) units. What value of x will make it a rhombus? 10 4 8 5

OpenStudy (anonymous):

@SourPandaBear @Callisto @ganeshie8

Parth (parthkohli):

Remember that a rhombus has all the sides EQUAL. We, therefore, have to make \(3x - 7 \) equal to \(x + 13\). An equation will help:\[3x - 7 = x + 13 \]Solve for \(x\).

OpenStudy (anonymous):

3x-7=x+13 -x -x 2x-7=13 +7 +7 2x=20 /2 /2 x=10

OpenStudy (anonymous):

FYI i may get in trouble but 10 is the answer